Using the digits 0,1,2,3,4,6 and 7. How many 3-digit even numbers between 200 and 600 can be made? What will be their sum?
Found 2 solutions by greenestamps, Edwin McCravy:
Answer by greenestamps(13200)



You didn't say whether digits could be repeated. I guessed that they could not. Of course, if they can be repeated, then my response is of no use to you.

Assuming digits can't be repeated....

(1) For the 3-digit number to be between 200 and 600, the first digit must be 2, 3, or 4.

(2) To be even, the last digit must be 2, 4, or 6; and it can't be the same as the first digit.

Together, those requirements give us the following list of 3-digit patterns, where the middle digit "_" can be any one of the remaining 5 digits:

2_4
2_6
3_2
3_4
3_6
4_2
4_6

That is 7 patterns, with 5 different values for the middle digit in each pattern; that makes a total of 35 3-digit numbers that can be made.

ANSWER to the first question: there are 35 3-digit numbers that can be made.

An analysis can be made to find the sum of those 35 numbers without simply listing them all out and adding them... but the straight addition might be faster and easier.

But let's think about how we can find the sum of all 35 numbers by looking at the sum of the digits in each column.

The first (hundreds) digit is 2 10 times, 3 15 times, and 4 10 times. The sum of the hundreds digits of all 35 numbers is then

(20+45+40)*100 = 10500

The digits in the ones column are 2 10 times, 4 10 times, and 6 15 times. The sum of the units digits of all 35 numbers is then

(20+40+90)*1 = 150

For the digits in the tens column, consider each of the 7 digits and see how many of the 7 patterns can have that digit as the middle digit.

0 can be used in all 7 patterns
1 can be used in all 7 patterns
2 can be used in 3 patterns
3 can be used in 4 patterns
4 can be used in 3 patterns
6 can be used in 4 patterns
7 can be used in all 7 patterns

The sum of the digits in the tens column of all 35 number is

((7*0)+(7*1)+(3*2)+(4*3)+(3*4)+(4*6)+(7*7))*10 = (0+7+6+12+12+24+49)*10 = 110*10 = 1100.

So the sum of all 35 of the numbers is

10500+1100+150 = 11750

ANSWER to the second question: the sum of the 35 numbers is 11750.

Answer by Edwin McCravy(20055)

The solution above is WRONG! 
Here's the correct solutions, for both inclusive and exclusive of 200 and 600:

Using the digits 0,1,2,3,4,6 and 7, how many 3-digit even numbers between
200 and 600 can be made?
You didn't state whether it was inclusive or exclusive of 200 and 600,
If it's 'exclusive' the smallest even number allowed is 202, and the largest
even number allowed is 576.  If 'inclusive' both 200 and 600 will be included.

However, for calculation purposes, it will be more convenient to include
200 and exclude 600. Then we can adjust the count or sum so as to
include or exclude 200 and 600.

So we will first assume that the smallest even number allowed is 200 and
the largest is 476. That is, we are including 200 but excluding 600.  We
will adjust later.

The hundreds digit can only be 2,3 or 4.
So that's 3 ways to choose the hundreds digit.
The tens digit can be any of the 7 digits 0,1,2,3,4,6 or 7
So that's 7 ways to choose the tens digit.
The ones digit must be 0,2,4 or 6, to make sure it's even.
So that's 4 ways to choose the ones digit.

So the answer we get here is 3•7•4 = 84 ways.
Now that includes the number 200, which we are excluding,
So we must subtract one from the 84.

Answer: 83, assuming we are not counting 200 or 600.
If we are including them both then the answer would be 2 more than 83, or 85.
----------------------------

What is their sum?
Think of the long column of 84 3-digit numbers to be added.  For
convenience, let's INclude 200 but EXclude 600.


  200
  202
  204
  206
  ...
  470
  472
  474
 +476
------
  sum


The 84 hundreds digits in the leftmost column contain an equal number of the 3
digits 2, 3 and 4. Since 84/3 = 28. the sum of all the hundreds digits is 28(2+3+4) = 252.
Therefore the sum of all hundreds digits contributes 252∙100 or 25200 to
the final sum.

The 84 tens digits in the middle column contain an equal number of all 7 digits
0,1,2,3,4,6 and 7.  Since 84/7 = 12. the sum of all the hundreds digits is 12(0+1+2+3+4+6+7) = 276.
Therefore the sum of all tens digits contributes 276∙10 or to the sum.

The 84 ones digits in the rightmost column contain an equal number of the 4
digits 0,2, 4 and 6. Since 84/4 = 21. the sum of all the hundreds digits is
21(0+2+4+6) = 252.

Therefore the sum of all ones digits contributes 252 to the final sum.

Therefore the sum which includes 200 but excludes 600 is 25200+2560+252 = 28212.
(Chances are this is NOT the answer you want, but it is the easiest to calculate
and from which the other two possible answers are easily calculated.]

If we want the sum which excludes both 200 and 600 we must subtract 200 from
28212 and get 28012.

If we want the sum which includes both 200 and 600 we must add 600 and get
28812.

Edwin